Bryan is twice as old as Colin. Colin is 4 years younger then Michael. In three years bryan Colin and Michael will be 2 more than half a century.how old are they now?
Answer by ankor@dixie-net.com(22740)

let b = Bryan's present age
let c = Colin's "
let m = Micheal's "
:
Write an equation for each statement:
:
"Bryan is twice as old as Colin."
b = 2c
:
"Colin is 4 years younger then Michael."
c = m - 4
or
m = c + 4
:
"In three years bryan Colin and Michael will be 2 more than half a century."
(b+3) + (c+3) + (m+3) = 52
b + c + m + 9 = 52
b + c + m = 52 - 9
b + c + m = 43
Replace b with 2c, replace m with (c+4)
2c + c + (c+4) = 43
4c = 43 - 4
4c = 39
c = 39/4
c = 9.75 yrs is Colin's present; usually expect an integer, but here it is
:
b = 2(9.75)
b = 19.5 yrs is Bryan's present age
:
m = 9.75 + 4
m = 13.75 yrs is Michael's present age